A payday loan is structured to obscure the true interest rate you are paying For example, in Washington, you pay a $30 fee
Answer #1

А. 1 Annual interest rate 2 Effective interest rate B 380.49% 3384.69%

calc:

A B 1 Annual interest rate =RATE(1,,-205,235)* 26 2 Effective interest rate =(1+B1/26)^26-1